nayira:

Find the expression that represents the length of the hypotenuse of a right triangle whose legs measure 2004-04-01-02-00_files/i0350000.jpgand 2004-04-01-02-00_files/i0350001.jpg. A. 2004-04-01-02-00_files/i0350002.jpg B. 2004-04-01-02-00_files/i0350003.jpg C. 2004-04-01-02-00_files/i0350004.jpg D. 2004-04-01-02-00_files/i0350005.jpg

dude:

Corrected Question: Find the expression that represents the length of the hypotenuse of a right triangle whose legs measure 2x\(^2\)-y\(^2\) and 2xy.

dude:

Okay use pythagoreans theorem Do you know what that equation is?

dude:

Welp you left Pythagorean Theorem a\(^2\)+b\(^2\)=c\(^2\) a and b are the legs, so we just substitute that into the equation (2x\(^2\)-y\(^2\))\(^2\)+(2xy)\(^2\)=c\(^2\) First simplify the legs (a and b) =\((2x^2-y^2)^2\) \(=4x^4-4x^2y^2+y^4\) =\((2xy)^2\) \(=4x^2y^2\) So we now have \((4x^4-4x^2y^2+y^4)+4x^2y^2\) Combine like terms \(4x^4\color{red}{-4x^2y^2}+y^4\color{red}{+4x^2y^2}\) =\(4x^4+y^4\) So we have \(4x^4+y^4=c^2\) Now we take the square root on both sides \(c=\sqrt{4x^4+y^4}\) Which happens to be A